0 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

A mechanism is given involving the nucleophilic attack of methanol (CH3OH)\left( {{\rm{C}}{{\rm{H}}_3}{\rm{OH}}} \right) on a sulfonyl chloride molecule (RSO2Cl)\left( {{\rm{RS}}{{\rm{O}}_2}{\rm{Cl}}} \right) ; R= phenyl. The cleavage whether it is C-O or O-H needs to be predicted. Follow the curved arrow mechanism to decide the cleavage.

Fundamentals

The mechanism of a reaction indicates the mode through which a reaction is taking place that is how the reactants are undergoing changes to form products.

The curved arrows represent the direction of electron flow. The tail of the curved arrow represents the donor atom from where attack is taking place. The head of the arrow represents the site at which attack is taking place that is the electron acceptor site. Generally, an arrow shows movement of electron pair thus making and breaking bond between atoms.

The reaction is as follows:

Io: ö
-R-30-
=n=
0:
HO
-R

The next step is as follows:

-I
:o=
I—:
-
R-
O
-
R-
S-
O:

The last step is as follows:

I
-
-
R-
co==O:
---
S-
0-
:0
-
Ho=n=0:
R-S-
0-
+ H2O

Ans:

OC-O bond cleavage
0-H bond cleavage
O Both C-O and O-H bond cleavage